# Unitarity, Causality, and Solar System Bounds, Significantly Limit Using Gravitational Waves to Test General Relativity

### v2

@@ 1 Introduction
Searches for either new effects or corrections that deviate from general relativity (GR) is an ongoing front in both the experimental and theoretical aspects of gravitation. On the experimental side, we have a century of precision tests of GR in the solar system. However, within the solar system, the spatial curvature is very small. More recently, we have LIGO/Virgo [LIGOScientific:2016aoc] detecting gravitational waves from merging black holes, where the spatial curvature is larger. Currently, all observations are broadly compatible with the predictions of GR. On the theoretical side, much has been devoted to understanding the space of possible deviations. Within effective field theory (EFT), the only way deviations can appear is through higher derivative operators and their possible appearance within an ultra-violated (UV) completion.
Constraining higher derivative operators from LIGO/Virgo observations has grown in interest. In particular, Ref. [Sennett:2019bpc] has put a bound on the quartic power of Riemann operator of $\Lambda_{4,bd}=1/(150\>\text{km})$. In a more EFT-bootstrap manner, [Caron-Huot:2021rmr, Caron-Huot:2022ugt] has constrained the cubic and quartic power of Riemann curvature operators to high numerical precision using constraints on the coefficients of the operators via causality. Whether or not these two regimes tell us something new about the Wilson coefficients of the higher curvature operators is an open question, and one we intend to begin answering.
@@ 2.1 Building the Effective Field Theory
The effective field theory action for gravity is an expansion in derivatives. We set $\hbar=c=1$ with metric signature (+,,,)(+,-,-,-)( + , - , - , - ) (and adapting the notation of [Caron-Huot:2022ugt]) we write the expansion as
@@
We have only included terms that involve powers of Riemann, but not terms that are powers of Ricci or Ricci tensor, such as $R^{2}$ or $R^{3}$, etc. The reason being is that we are primarily interested in the merger of Kerr black holes. Such black holes have $R=0$ and $R_{\mu\nu}=0$ as they are vacuum solutions. So we only need to focus on powers of Riemann, as this is non-zero even for vacuum solutions. One could consider these as local terms, $R^{2},\,R_{\mu\nu}R^{\mu\nu}$, etc, and enquire about their possible corrections to Newton’s law. We comment on this issue in Appendix C.
@@ 2.2 Constraints from Observed Gravitational Waves
The interesting leading order terms start at cubic and quartic order in the Riemann tensor. In Ref. [Sennett:2019bpc], bounds are placed on these operators from comparison to LIGO/Virgo data (although the cubic term is ignored, for reasons we will come to below). We can express these bounds in terms of a physical scales (inverse lengths) $\Lambda_{3}$ and $\Lambda_{4}$ that sets the size of the cubic and quartic powers of the Riemann operator in the action. These are related to our Wilson coefficients as
@@
The observational bound found in [Sennett:2019bpc] is
@@
One might wonder if at this scale the theory undergoes strong coupling and changes radically at the scale $L\sim 1/\Lambda_{3,4}$ (which would need to be close to $\sim 150\,\text{km}$ to have relevance to existing observations) to be replaced by some new physics. However, since we have tested gravity to much smaller scales, which has produced nothing new, this seems to be a big problem. In response to this, in [Sennett:2019bpc] it is suggested that below this scale there is a “soft UV completion” in which the new physics stays hidden and Einstein gravity remains intact. While this is a priori possible, in this work we would like to test this possibility against a concrete UV completion, as we come to shortly.
@@ 2.3 Constraints from Causality
As has been known for the past few years [Camanho:2014apa], the inclusion of $(\text{Riem})^{3}$ in any effective action for gravitation produces superluminality in a Lorentz invariant theory and therefore acausality. However, as was shown in [Bern:2021ppb] and [Caron-Huot:2022ugt], when the quartic operators are included, then non-zero cubic terms are allowed, so long as an inequality between the cubic and quartic coefficients (which will be explained more in-depth in section 2.4), is satisfied. The inequality is
@@
where $m_{\text{min}}$ is the mass-gap associated with the mass of the lightest particle that has been integrated out and $\delta$ is an $\mathcal{O}(1)$ prefactor. This inequality shows that it is not possible to turn on the cubic coupling without including the quartic coupling also. This inequality was originally conjectured and argued for in Ref. [Camanho:2014apa] where it was suggested $\delta=0.25$, and confirmed by Ref. [Caron-Huot:2021rmr] where the more precise value is said to be $\delta=0.26$, and also by Ref. [Bern:2021ppb]. Furthermore, as is well known, the quartic coefficients need to obey positivity bounds, namely [^3] Let us also note that the cubic and quartic coefficients cannot be arbitrarily large. As argued in [Caron-Huot:2021rmr], they are bounded by $(\alpha_{3}^{2}+\tilde{\alpha}_{3}^{2})m_{\text{min}}^{8}\leq 24.9\log(m_{\text{min}}/m_{\text{IR}})-27.6$ (6) $(\alpha_{4}+\alpha_{4}^{\prime})m_{\text{min}}^{6}\leq 12.3\log(m_{\text{min}}/m_{\text{IR}})-13.5$ (7) where $m_{\text{IR}}$ is an infrared scale.
@@ 2.4 Higher Derivative Operators from Integrating Out Massive Particles
One can integrate out these scalars to build the low energy EFT organized in a derivative expansion. Specifically, the main method is via helicity amplitudes that compute graviton-graviton scattering in the EFT. At one-loop the coefficients of $(\text{Riem})^{3}$ and $(\text{Riem})^{4}$ have been determined [Bern:2021ppb], with values
$\alpha_{3}$ $=\sum_{i}\frac{1}{(4\pi)^{2}}\frac{1}{5040}\frac{1}{m_{i}^{2}}\left(\frac{\kappa^{2}}{2}\right),$ $\alpha_{4}=\sum_{i}\frac{1}{(4\pi)^{2}}\frac{11}{75600}\frac{1}{m_{i}^{4}}\left(\frac{\kappa}{2}\right)^{2},$ $\alpha_{4}^{\prime}=\sum_{i}\frac{1}{(4\pi)^{2}}\frac{1}{75600}\frac{1}{m_{i}^{4}}\left(\frac{\kappa}{2}\right)^{2}.$ (10)
@@
where $m_{\text{pl}}=1/\sqrt{G}$ is the Planck mass. This is a similar relation found previously in [Dvali:2007hz] relating particle species, their mass, and an EFT cutoff. By rewriting this bound in terms of the scales $\Lambda_{4}$ or $\Lambda_{3}$, this becomes the following pair of inequalities
@@
We can invert this logic and use it to bound $m_{\text{min}}$. The best case scenario for phenomenological purposes then is that all the masses are equal $m=m_{\text{min}}=m_{1}=\ldots=m_{N}$. If we use the observational bound on the EFT of Ref. [Sennett:2019bpc] of $\Lambda_{4,0}\simeq(150\text{km})^{-1}$ as a useful reference, then we obtain a bound on the mass of the lightest particle of
@@
with equality holding when we saturate the theoretical bound $m=m_{c}$. Although this is an extremely large $N$, it is important to note that such species are assumed to be in a hidden sector and not directly coupled to the Standard Model at tree level. Their only direct couplings to the Standard Model occur via quantum loops as we now explore and learn their possible observational consequences. One might expect that such a hige number of particles would already have been easily visible at colliders, for example from $e^{+}e^{-}$ annihilation into these scalars via the graviton in the s-channel; but it is important to recall that the proposal that is presented in Ref. [Sennett:2019bpc] is that of the “soft UV completion", meaning that one should not use the theory at arbitrarily low scales.
@@ 3.1 Massive Scalar Amplitude
In the presence of many light scalars, there can be sizable corrections to the gravitational potential between massive objects. The setup for computing such corrections to Newton’s potential was given in [Donoghue:1994dn] initially with the finalized 1-loop correction in pure gravity given in [Bjerrum-Bohr:2002gqz]. We give the basic setup for the EFT treatment of gravity and leave the further details to the references.
@@
where $\eta_{\mu\nu}$ is the Minkowski metric and $h_{\mu\nu}$ is the graviton field. By inserting into the Einstein-Hilbert action with the scalars and matter (9), one can systematically derive the Feynman rules we will be using in this paper; we summarize these in Appendix A.1. For brevity, we do not include the ghost diagrams from the path integral quantization of this gauge theory, as they will not play a direct role to the order we are working. The ghost diagram contributions have previously been computed in [Bjerrum-Bohr:2002gqz].
The setup for calculating the 1-loop correction to the potential between two point particles exchanging a graviton with a massive scalar at 1-loop is relatively straightforward. The difficulty is solely on the number of terms that arise, which is where we utilize Mathematica [Mathematica] and specifically the packages FeynCalc, Package-X, and FeynGrav [Mertig:1990an, Shtabovenko:2020gxv, Shtabovenko:2016sxi, Latosh:2023zsi]. We utilize dimensional regularization with the convention $d=4-2\epsilon$, where $d$ is the number of spacetime dimensions (we then take $\epsilon\to 0$ or $d\to 4$ at the end of the calculation). We also use Poincare symmetry to simplify integrals with integrands proportional to $p^{\mu}p^{\nu}p^{\alpha}p^{\beta}$ and $p^{\mu}p^{\nu}$ to integrals with integrands proportional to $p^{4}$ and $p^{2}$ times factors of the Minkowski metric $\eta^{\mu\nu}$.
@@
There are four diagrams we need to take into account which can be seen in equations (21) and (22) while the third and fourth will be discussed later. The general structure of the loop amplitude is found in equation (23) with the respective components of the matrix elements for equations (21) and (22) are given in equations (24) and (25) respectively.
@@
where $\Delta=m^{2}-x(1-x)q^{2}$ and $\mu$ is the scale at which we define the couplings, and we have suppressed the $+i0^{+}$ in the Feynman propagator for the sake of brevity. The 1PI changes for equation (22) as
@@
where the 4-point vertex between two gravitons and two scalars is given in Appendix A, equation (A6). We follow the typical procedure for handling these types of diagrams, utilizing the $(AB)^{-1}$ Feynman trick, and so on. The numerous amount of terms after contracting each 1PI with the external legs can be sorted into Passarino-Veltman coefficients that can in turn be evaluated from known integrals [Passarino:1978jh]. In the small $\epsilon$-expansion of the integrals with $d=4-2\epsilon$, we keep the divergent $\mathcal{O}(\epsilon^{-1})$ terms and the finite $\mathcal{O}(1)$ terms.
@@ 3.1.1 One-Loop Diagrams
Equation (22) has a relatively simple tensor structure; we report on that first. The Feynman diagram of equation (22) is given as
@@
where $A_{0}(m^{2})$ is the tadpole loop integral Passarino-Veltman coefficient [Passarino:1978jh],
@@
where the tensors $F^{\mu\nu\alpha\beta}$ and $G^{\mu\nu\alpha\beta}$ are defined in Appendix A. The evaluation of the Passarino-Veltman coefficients leads to the solution provided in [Donoghue:2022chi] but only up to $\mathcal{O}(m^{4})$. Here, we confirm this result, and include the $\mathcal{O}(m^{2})$ and $\mathcal{O}(m^{0})$ terms as well. The result is given in the Appendix in Eq. (A11).
We now contract with the external legs as defined in the appendix (A4). In doing so, we also take the NR limit as described in equation (26). The diagram of (22) becomes
@@
$+\frac{i\kappa^{4}M_{1}^{2}M_{2}^{2}}{115200\pi^{2}\mathbf{q}^{6}}\biggm{(}\!15\left(131+30\log\left(\frac{4\pi\mu^{2}e^{-\gamma}}{m^{2}}\right)\right)m^{4}\mathbf{q}^{2}+10\left(49+15\log\left(\frac{4\pi\mu^{2}e^{-\gamma}}{m^{2}}\right)\right)m^{2}\mathbf{q}^{4}$
$+\left(-12+45\log\left(\frac{4\pi\mu^{2}e^{-\gamma}}{m^{2}}\right)\right)\mathbf{q}^{6}$

### v3

@@ 1 Introduction
Searches for either new effects or corrections that deviate from general relativity (GR) is an ongoing front in both the experimental and theoretical aspects of gravitation. On the experimental side, we have a century of precision tests of GR in the solar system. However, within the solar system, the spatial curvature is very small. More recently, we have LIGO/Virgo LIGOScientific:2016aoc detecting gravitational waves from merging black holes, where the spatial curvature is larger. Currently, all observations are broadly compatible with the predictions of GR. On the theoretical side, much has been devoted to understanding the space of possible deviations. Within effective field theory (EFT), the only way deviations can appear is through higher derivative operators and their possible appearance within an ultra-violated (UV) completion.
Constraining higher derivative operators from LIGO/Virgo observations has grown in interest. In particular, Ref. Sennett:2019bpc has put a bound on the quartic power of Riemann operator of $\Lambda_{4,bd}=1/(150\>\text{km})$. In a more EFT-bootstrap manner, Caron-Huot:2021rmr ; Caron-Huot:2022ugt has constrained the cubic and quartic power of Riemann curvature operators to high numerical precision using constraints on the coefficients of the operators via causality. Whether or not these two regimes tell us something new about the Wilson coefficients of the higher curvature operators is an open question, and one we intend to begin answering. An analysis of corrections to Newton’s potential from parity-preserving operators has been considered in Melville:2024zjq with time delay implications.
@@ 2.1 Building the Effective Field Theory
The effective field theory action for gravity is an expansion in derivatives. We set $\hbar=c=1$ with metric signature (+,,,)(+,-,-,-)( + , - , - , - ) (and adapting the notation of Caron-Huot:2022ugt ) we write the expansion as
@@
We have only included terms that involve powers of Riemann, but not terms that are powers of Ricci or Ricci tensor, such as $R^{2}$ or $R^{3}$, etc. The reason being is that we are primarily interested in the merger of Kerr black holes. Such black holes have $R=0$ and $R_{\mu\nu}=0$ as they are vacuum solutions. So we only need to focus on powers of Riemann, as this is non-zero even for vacuum solutions. One could consider these as local terms, $R^{2},\,R_{\mu\nu}R^{\mu\nu}$, etc, and inquire about their possible corrections to Newton’s law. We comment on this issue in Appendix C.
@@
Hence, the interesting leading order terms start at cubic and quartic order in the Riemann tensor.
@@ 2.2 Constraints from Observed Gravitational Waves
In Ref. Sennett:2019bpc , bounds are placed on these operators from comparison to LIGO/Virgo data (although in Ref. Sennett:2019bpc the cubic term was assumed small, for reasons we will come to below). We can express these bounds in terms of a physical scales (inverse lengths) $\Lambda_{3}$ and $\Lambda_{4}$ that sets the size of the cubic and quartic powers of the Riemann operator in the action. These are related to our Wilson coefficients as
@@
The observational bound found in Sennett:2019bpc is
@@
One might wonder if at this scale the theory undergoes strong coupling and changes radically at the scale $L\sim 1/\Lambda_{3,4}$ (which would need to be close to $\sim 150\,\text{km}$ to have relevance to existing observations) to be replaced by some new physics. However, since we have tested gravity to much smaller scales, which has produced nothing new, this seems to be a big problem. In response to this, in Sennett:2019bpc it is suggested that below this scale there is a “soft UV completion” in which the new physics stays hidden and Einstein gravity remains intact. While this is a priori possible, in this work we would like to test this possibility against a concrete UV completion, as we come to shortly.
@@ 2.3 Constraints from Causality
As has been known for the past few years Camanho:2014apa , if one only includes $(\text{Riem})^{3}$ in the effective action for gravitation, and no other higher dimension operators, then there is superluminality. However, when the quartic operators $(\text{Riem})^{4}$ are included, then non-zero cubic terms are allowed, so long as an inequality between the cubic and quartic coefficients (which will be explained more in-depth in section 2.4), is satisfied (see Bern:2021ppb and Caron-Huot:2022ugt ). The inequality is
@@
where $m_{\text{min}}$ is the mass-gap associated with the mass of the lightest particle that has been integrated out and $\delta$ is an $\mathcal{O}(1)$ prefactor. This inequality shows that it is not possible to turn on the cubic coupling without including the quartic coupling also. This inequality was originally conjectured and argued for in Ref. Camanho:2014apa where it was suggested $\delta=0.25$, and confirmed by Ref. Caron-Huot:2021rmr where the more precise value is said to be $\delta=0.26$, and also by Ref. Bern:2021ppb . Furthermore, as is well known, the quartic coefficients need to obey positivity bounds, namely [^1] Let us also note that the cubic and quartic coefficients cannot be arbitrarily large. As argued in Caron-Huot:2021rmr , they are bounded by $(\alpha_{3}^{2}+\tilde{\alpha}_{3}^{2})m_{\text{min}}^{8}\leq 24.9\log(m_{\text{min}}/m_{\text{IR}})-27.6$ (6) $(\alpha_{4}+\alpha_{4}^{\prime})m_{\text{min}}^{6}\leq 12.3\log(m_{\text{min}}/m_{\text{IR}})-13.5$ (7) where $m_{\text{IR}}$ is an infrared scale.
@@ 2.4 Higher Derivative Operators from Integrating Out Massive Particles
One can integrate out these scalars to build the low energy EFT organized in a derivative expansion. Specifically, the main method is via helicity amplitudes that compute graviton-graviton scattering in the EFT. At one-loop the coefficients of $(\text{Riem})^{3}$ and $(\text{Riem})^{4}$ have been determined Bern:2021ppb , with values
$\alpha_{3}=\sum_{i}\frac{1}{(4\pi)^{2}}\frac{1}{5040}\frac{1}{m_{i}^{2}}\left(\frac{\kappa^{2}}{2}\right),$ $\>\>\>\alpha_{4}=\sum_{i}\frac{1}{(4\pi)^{2}}\frac{11}{75600}\frac{1}{m_{i}^{4}}\left(\frac{\kappa}{2}\right)^{2},$
$\alpha_{4}^{\prime}=\sum_{i}\frac{1}{(4\pi)^{2}}$ $\frac{1}{75600}\frac{1}{m_{i}^{4}}\left(\frac{\kappa}{2}\right)^{2}.$ (10)
@@
where $m_{\text{pl}}=1/\sqrt{G}$ is the Planck mass. This is a similar relation found previously in Dvali:2007hz relating particle species, their mass, and an EFT cutoff. By rewriting this bound in terms of the scales $\Lambda_{4}$ or $\Lambda_{3}$, this becomes the following pair of inequalities
@@
We can invert this logic and use it to bound $m_{\text{min}}$. The best case scenario for phenomenological purposes then is that all the masses are equal $m=m_{\text{min}}=m_{1}=\ldots=m_{N}$. If we use the observational bound on the EFT of Ref. Sennett:2019bpc of $\Lambda_{4,0}\simeq(150\text{km})^{-1}$ as a useful reference, then we obtain a bound on the mass of the lightest particle of
@@
with equality holding when we saturate the theoretical bound $m=m_{c}$. Although this is an extremely large $N$, it is important to note that such species are assumed to be in a hidden sector and not directly coupled to the Standard Model at tree level. Their only direct couplings to the Standard Model occur via quantum loops as we now explore and learn their possible observational consequences. One might expect that such a huge number of particles would already have been easily visible at colliders, for example from $e^{+}e^{-}$ annihilation into these scalars via the graviton in the s-channel; but it is important to recall that the proposal that is presented in Ref. Sennett:2019bpc is that of the “soft UV completion", meaning that one should not use the theory at arbitrarily low scales.
@@ 3.1 Massive Scalar Amplitude
In the presence of many light scalars, there can be sizable corrections to the gravitational potential between massive objects. The setup for computing such corrections to Newton’s potential was given in Donoghue:1994dn initially with the finalized 1-loop correction in pure gravity given in Bjerrum-Bohr:2002gqz . We give the basic setup for the EFT treatment of gravity and leave the further details to the references.
There have been many other studies on loop corrections to physical observables such as time-delay effects from violations of causality for example in Hollowood:2015elj ; deRham:2020zyh ; Chen:2021bvg ; Caron-Huot:2022jli ; deRham:2022gfe ; Bhat:2023puy ; Caron-Huot:2024tsk ; AccettulliHuber:2020oou ; AccettulliHuber:2020dal . However, our focus is on corrections to Newton’s potential. And furthermore, we shall determine its observational consequences for the solar system.
@@
where $\eta_{\mu\nu}$ is the Minkowski metric and $h_{\mu\nu}$ is the graviton field. By inserting into the Einstein-Hilbert action with the scalars and matter (9), one can systematically derive the Feynman rules we will be using in this paper; we summarize these in Appendix A.1. For brevity, we do not include the ghost diagrams from the path integral quantization of this gauge theory, as they will not play a direct role to the order we are working. The ghost diagram contributions have previously been computed in Bjerrum-Bohr:2002gqz .
The setup for calculating the 1-loop correction to the potential between two point particles exchanging a graviton with a massive scalar at 1-loop is relatively straightforward. The difficulty is solely on the number of terms that arise, which is where we utilize Mathematica Mathematica and specifically the packages FeynCalc, Package-X, and FeynGrav Mertig:1990an ; Shtabovenko:2020gxv ; Shtabovenko:2016sxi ; Latosh:2023zsi . We utilize dimensional regularization with the convention $d=4-2\epsilon$, where $d$ is the number of spacetime dimensions (we then take $\epsilon\to 0$ or $d\to 4$ at the end of the calculation). We also use Poincare symmetry to simplify integrals with integrands proportional to $p^{\mu}p^{\nu}p^{\alpha}p^{\beta}$ and $p^{\mu}p^{\nu}$ to integrals with integrands proportional to $p^{4}$ and $p^{2}$ times factors of the Minkowski metric $\eta^{\mu\nu}$.
@@
There are four diagrams we need to take into account which can be seen in equations (21) and (3.1) while the third and fourth will be discussed later. The general structure of the loop amplitude is found in equation (22) with the respective components of the matrix elements for equations (21) and (3.1) are given in equations (23) and (24) respectively.
@@
where $\Delta=m^{2}-x(1-x)q^{2}$ and $\mu$ is the scale at which we define the couplings, and we have suppressed the $+i0^{+}$ in the Feynman propagator for the sake of brevity. The 1PI changes for equation (3.1) as
@@
where the 4-point vertex between two gravitons and two scalars is given in Appendix A, equation (A6). We follow the typical procedure for handling these types of diagrams, utilizing the $(AB)^{-1}$ Feynman trick, and so on. The numerous amount of terms after contracting each 1PI with the external legs can be sorted into Passarino-Veltman coefficients that can in turn be evaluated from known integrals Passarino:1978jh . In the small $\epsilon$-expansion of the integrals with $d=4-2\epsilon$, we keep the divergent $\mathcal{O}(\epsilon^{-1})$ terms and the finite $\mathcal{O}(1)$ terms.
@@ 3.1.1 One-Loop Diagrams
Equation (3.1) has a relatively simple tensor structure; we report on that first. The Feynman diagram of equation (3.1) is given as
@@
where $A_{0}(m^{2})$ is the tadpole loop integral Passarino-Veltman coefficient Passarino:1978jh ,
@@
where the tensors $F^{\mu\nu\alpha\beta}$ and $G^{\mu\nu\alpha\beta}$ are defined in Appendix A. The evaluation of the Passarino-Veltman coefficients leads to the solution provided in Donoghue:2022chi but only up to $\mathcal{O}(m^{4})$. Here, we confirm this result, and include the $\mathcal{O}(m^{2})$ and $\mathcal{O}(m^{0})$ terms as well. The result is given in the Appendix in Eq. (A11).
We now contract with the external legs as defined in the appendix (A4). In doing so, we also take the NR limit as described in equation (25). The diagram of (3.1) becomes
@@
$+\frac{i\kappa^{4}M_{1}^{2}M_{2}^{2}}{115200\pi^{2}\mathbf{q}^{6}}\biggm{(}\!15\left(131+30\log\left(\frac{4\pi\mu^{2}e^{-\gamma}}{m^{2}}\right)\right)m^{4}\mathbf{q}^{2}$
$+10\left(49+15\log\left(\frac{4\pi\mu^{2}e^{-\gamma}}{m^{2}}\right)\right)m^{2}\mathbf{q}^{4}+\left(-12+45\log\left(\frac{4\pi\mu^{2}e^{-\gamma}}{m^{2}}\right)\right)\mathbf{q}^{6}$